OpenStudy (anonymous):

Tracy needs to find the length and width of the equipment storage room where she works. The room is 1 meter longer than it is wide. Its perimeter is 14 meters. Find the width in meters

OpenStudy (anonymous):

let the width = x meter then lenght = x+1meter then parameter is x +(x+1)+x+(x+1) = 14 now solve for x

OpenStudy (anonymous):

4x + 2? Dunno how to do this, I'm horrible in Algebra lol!

OpenStudy (anonymous):

actually the room in the shape of rectangle the parameter of a rectangle is sum of the length of all its sides as here the width is x and length is x+1 now the rectangle has four sides thus parameter is = x+x+(x+1)+(x+1) = 4x+2

OpenStudy (anonymous):

Ok so I was right about the 4x + 2 thing?

OpenStudy (texaschic101):

The perimeter of a rectangle is : P = 2L + 2W P = 14 L = W + 1 Now sub your stuff into the formula

OpenStudy (texaschic101):

14 = 2(W + 1) + 2W 14 = 2W + 2 + 2W can you take it from here and solve for W

OpenStudy (anonymous):

Well, do you combine like terms? If not, then no. I know this involves dividing one side by something else to get W by it's self, correct?

OpenStudy (texaschic101):

combine like terms

OpenStudy (mathstudent55):

@Skiller8860 You are correct. The left side of the equation becomes 4x + 2 That means the equation now is 4x + 2 = 14 Now you need to solve the two-step equation 4x + 2 = 14 First, subtract 2 from both sides to get 4x by itself on the left side. Then divide both sides by 4 to find x.

OpenStudy (anonymous):

3 + 3W when combining like terms, correct? Or do you do this, 2 * W = 2W 2 * 1 = 2 Therefore it equals, 2W + 2 + 2W So that means 4W + 2 correct @texaschic101

OpenStudy (anonymous):

Which one is correct, first method or last and @mathstudent55 X = 3 then?

OpenStudy (texaschic101):

yes...14 = 4W + 2 -- subtract 2 from both sides 14 - 2 = 4W 12 = 4W -- divide by 4 12/4 = W 3 = W -- width is 3 L = W + 1 L = 3 + 1 L = 4 -- length is 4 Mathstudent...what did I do wrong here...because when I check it back into the formula, it does not equal 14

OpenStudy (texaschic101):

oh...wait a sec P = 2L + 2W 14 = 2(4) + 2(3) 14 = 8 + 6 14 = 14 (correct) nvm mathstudent......I just messed up in the checking....it checks out correct L = 4 and W = 3

OpenStudy (anonymous):

So the answer is 3? =]

OpenStudy (texaschic101):

yes...W = 3 meters

OpenStudy (anonymous):

Thank you so much!!!!!!!! =]

OpenStudy (texaschic101):

do you have any questions Skiller ?

OpenStudy (texaschic101):

glad to help :)

OpenStudy (anonymous):

Nope, thanks :D. Wrote you a testimonial haha =].